OpenStudy (anonymous):

Triq Identities Anybody. find tan theta if sin theta=1/4

OpenStudy (anonymous):

you cannot but you can narrow it down a bit

OpenStudy (anonymous):

ok

OpenStudy (anonymous):



OpenStudy (anonymous):

there is a picture of an angle \(\theta\) where \(\sin(\theta)=\frac{1}{4}\) what is missing is the third side of the triangle, which you find via pythagoras

OpenStudy (anonymous):

oh the pythagoream theorem...which i spelled wrong but thanks alot

OpenStudy (anonymous):

also

OpenStudy (anonymous):

find cos theta if sin theta=1/4 and falls in quadrant 2

OpenStudy (jdoe0001):

 so you'd use the pythagorean theorem again now, we're in the 2nd Quadrant, keep in mind that one the 2nd Quadrant sine is positive, cosine is negative the pythagorean theorem doesn't give a + or - root, just a root the Quadrant the angle is in, narrows it down